OpenStudy (anonymous):

What is the solution set for |2x – 3| = 17?

OpenStudy (anonymous):

2x-3=17 2x=20 x=10 -2x+6=17 -2x=11 x=-11/2

OpenStudy (anonymous):

x=-7 x=10

OpenStudy (anonymous):

For the second part of atjari: -2x+3=17 -2x=14 x=-7

OpenStudy (anonymous):

thanks you guys

OpenStudy (anonymous):

Ya u r right Konsaliaris.